OpenStudy (ja1):

What are the solutions for:

OpenStudy (ja1):

What are the solutions for x in the proportion \[\frac{ 2x + 3 }{ x } = \frac{ 2x }{ 2 }\]

OpenStudy (anonymous):

2 x + 3/x = x

OpenStudy (ja1):

Ok sorry I was AFK

OpenStudy (ja1):

@sammaraq60 I am here to understand how to get to the answer not to get an answer.

OpenStudy (ja1):

So please continue but help me understand how to do this.

OpenStudy (luigi0210):

Cross multiply first, do you know how to do that? \

OpenStudy (anonymous):

write the left hand side as a single fraction

OpenStudy (ja1):

Yes but I think I am doing something wrong because my answer is wrong.

OpenStudy (anonymous):

bring x + 3/x together using the common denominator

OpenStudy (luigi0210):

Well cross multiply and show me what you get

OpenStudy (anonymous):

it should look like this x^2 +3/x =0

OpenStudy (ja1):

But don't I need to first simplify what I ca to do this?

OpenStudy (ja1):

How did you take away the 2 from 2x + 3 samm?

OpenStudy (anonymous):

before you crossmultiply , simplify 2x/2= x

OpenStudy (ja1):

Ok gotcha so I would just mutiply the denominator and the denominator by x?

OpenStudy (anonymous):